# Class-RAG: Content Moderation with Retrieval Augmented Generation

### v1

@@ Abstract
Robust content moderation classifiers are essential for the safety of Generative AI systems. Content moderation, or safety classification, is notoriously ambiguous: differences between safe and unsafe inputs are often extremely subtle, making it difficult for classifiers (and indeed, even humans) to properly distinguish violating vs. benign samples without further context or explanation. Furthermore, as these technologies are deployed across various applications and audiences, scaling risk discovery and mitigation through continuous model fine-tuning becomes increasingly challenging and costly. To address these challenges, we propose a Classification approach employing Retrieval-Augmented Generation (Class-RAG). Class-RAG extends the capability of its base LLM through access to a retrieval library which can be dynamically updated to enable semantic hotfixing for immediate, flexible risk mitigation. Compared to traditional fine-tuned models, Class-RAG demonstrates flexibility and transparency in decision-making. As evidenced by empirical studies, Class-RAG outperforms on classification and is more robust against adversarial attack. Besides, our findings suggest that Class-RAG performance scales with retrieval library size, indicating that increasing the library size is a viable and low-cost approach to improve content moderation.
@@ 1 Introduction
Recent advances in Generative AI technology have enabled new generations of product applications, such as text generation OpenAI (2023); Anthropic (2023); Dubey (2024), text-to-image generation Ramesh et al. (2021); Dai et al. (2023); Rombach et al. (2022), and text-to-video generation Meta (2024). Consequently, the pace of model development must be matched by the development of safety systems which are properly equipped to mitigate novel harms, ensuring the system’s overall integrity. This is important to prevent the use of Generative AI products from being exploited by bad actors to disseminate misinformation, glorify violence, and proliferate sexual content Foundation (2023).
Traditional model fine-tuning approaches are often employed to this end, and classifiers learning patterns from labeled content moderation text data are leveraged as guardrails for these deployed systems OpenAI (2023). However, there are many challenges associated with automating the content moderation task with a traditional model fine-tuning approach. First, content moderation is a highly subjective task, meaning that inter-annotator agreement in labeled data is low, due to, potentially, many different interpretations of policy guidelines, especially on borderline cases Markov et al. (2023). Second, it is impossible to enforce a universal taxonomy of harm, not only due to the subjectivity of the task, but due to the impact of systems scaling to new locales, new audiences, and new use cases, with different guidelines and different gradients of harm defined on those guidelines Shen et al. (2024). Training a robust content moderation classifier is already challenging due to the subjective nature of the task and the limitations of structured labeled data in capturing this subjectivity. Additional challenges arise when classifiers must be continuously fine-tuned to adapt to an evolving landscape of safety risks.
To address these challenges of subjectivity and inflexibility as a result of scale, we propose a Classification approach to content moderation which employs Retrieval-Augmented Generation (Class-RAG) to add context to elicit reasoning for content classification. While RAG Lewis et al. (2020) is often used for knowledge-intensive tasks, we find it to be useful for ambiguous tasks as well. Our content moderation system consists of an embedding model, a retrieval library consisting of both negative and positive examples, a retrieval module, and a fine-tuned LLM classifier. When a user inputs a query, we retrieve the most similar negative and positive examples, and enrich the original input query to the classifier with the contextual information derived from similar retrieved queries.
Main contributions
@@
  • Improved Classification Performance On experiments, Class-RAG demonstrates superior classification performance compared to either fine-tuning a lightweight 4-layer Transformer pre-trained on content moderation data or fine-tuning a general-purpose 8b parameter LLM.
  • Increased Flexibility By having an easily updated retrieval library, Class-RAG enables low-cost customization for various applications, seamlessly adapting to policy changes without requiring model retraining. The consequent ability to enable semantic (instead of hard) blocking allows for more precise identification and filtering of sensitive content.
  • Scalability and Cost-Effectiveness Our findings indicate that Class-RAG’s performance scales with the size of the retrieval library, suggesting that increasing the library size presents a viable and cost-effective approach to enhancing classification performance.
@@ 2 Related Work
Content moderation
Much work has been done in the last decade to mitigate the dissemination of undesired content in the wake of innovations in communication technologies. Machine learning approaches have been proposed to address sentiment classification Yu et al. (2017), harassment Yin et al. (2009), hate speech detection Gambäck and Sikdar (2017), abusive language Nobata et al. (2016), and toxicity Adams et al. (2017). General improvements in deep learning have also accelerated the field of content moderation. WPIE, or Whole Post Integrity Embeddings, is built with BERT and XLM on top of advances in self-supervision, and obtains a holistic understanding of a post through a pretrained universal representation of content Schroepfer (2019). Even more recent advances in Generative AI have also spurred the question of whether or not LLMs could potentially be used as content moderators Huang (2024).
Generative AI safety
With the advent of Generative AI also comes a proliferation of harm types beyond hate speech or toxicity detection whose mitigations and benchmarks warrant further research and exploration. A comprehensive AI harm taxonomy encompasses such harm categories like academic dishonesty, unauthorized privacy violations, and non-consensual nudity Zeng et al. (2024). As the approaches to mitigating Generative AI safety proliferate, so too do benchmarks which establish baselines for the efficacy of existing classifiers, such as UnsafeBench Qu et al. (2024). Benchmarks such as I2P  Schramowski et al. (2023a), which includes approximately 4.5k real-world examples of English text-to-image prompts, and P4D Chin et al. (2024), which employs red-teaming strategies for automated detection of unsafe outputs, have provided valuable datasets to evaluate harmful categories like self-harm, illegal activity, sexual content, and graphic violence.
@@
While the base capabilities of LLMs OpenAI (2023); Anthropic (2023); Dubey (2024) are powerful, LLMs have a known tendency to hallucinate Huang et al. (2023), a lack of ability to provide interpretable explanations for their generations, and are constrained by the cutoff date of their training data. Retrieval Augmented Generation (RAG) Lewis et al. (2020) mitigates some of these problems by augmenting the base capabilities of large pre-trained language models with a retrieval mechanism to explicit non-parametric memory Zhao et al. (2024). RAG has often been used to extend the capabilities of LLMs, particularly for knowledge-intensive tasks Gao et al. (2024). The flexibility of RAG-based approaches allows for applications that do not require additional in-domain finetuning, such as adapting detector models in the field of computer vision Jian et al. (2024). For example, RAFT improves the model’s ability to answer questions in open-book in-domain settings Zhang et al. (2024).
@@ 3 System Architecture
Figure 1: Diagram flow of Class-RAG
Class-RAG is a quadripartite system consisting of an embedding model, a retrieval library, a retrieval module, and a fine-tuned LLM classifier. When a user inputs a prompt, an embedding is computed on the prompt via the embedding model. The embedding of the user input is compared against an index of embeddings which are mapped to positive and negative examples of prompts in the retrieval library. Using Faiss, a library for efficient similarity search Douze et al. (2024), $k$ nearest reference examples are retrieved against the embedding of the user input prompt, and the reference examples and input prompt are then sent to the fine-tuned LLM for classification. We leverage the CoPro dataset Liu et al. (2024) to train and evaluate our model.
@@ 3.1 Embedding Model
We leverage the DRAGON RoBERTa Lin et al. (2023) context model as our primary embedding model. DRAGON is a bi-encoder dense retrieval model that utilizes a dual-encoder architecture to embed both queries and documents into dense vector representations, facilitating efficient retrieval of relevant information. In this study, we specifically employ the context encoder component of the DRAGON model. To investigate the impact of alternative embedding models on our approach, we also evaluate a variant of WPIE (Whole Post Integrity Embedding) Meta (2021). The WPIE model we test is a 4-layer XLM-R Conneau et al. (2020) model that has been pre-trained on content moderation data, yielding two distinct outputs: an unsafe probability estimation and a prompt embedding representation.
@@ 3.2 Retrieval Library
Our retrieval library is comprised of two distinct sub-libraries: a safe library and an unsafe library. Each entry in the retrieval library is represented by a quadruplet of attributes, including: (1) prompt, (2) label, (3) embedding, and (4) explanation. The construction of the retrieval library is described in detail in the Data Preparation section.
@@ 3.2.1 Retrieval Module
Given the selected embedding, we leverage the Faiss library for similarity search Douze et al. (2024) to efficiently retrieve the two nearest safe examples and the two nearest unsafe examples from the retrieval library. Specifically, we utilize the L2 distance metric to compute the similarity between the input embedding and the embeddings stored in the retrieval library, allowing us to identify the most relevant examples.
@@ 3.3 LLM Classifier
Inspired by LlamaGuard Inan et al. (2023), the classifier is fine-tuned on top of the OSS Llama-3-8b checkpoint Dubey (2024).
@@ 4.1 Dataset Details
We leverage the CoPro dataset Liu et al. (2024) to train and evaluate our model. The CoPro dataset consists of an in-distribution (ID) test set and an out-of-distribution (OOD) test set, both generated by a large language model. The OOD test set is generated with unseen seed inputs, providing a more challenging evaluation scenario. In addition to CoPro, we use the Unsafe Diffusion (UD) Qu et al. (2023) and I2P++ Liu et al. (2024) datasets to evaluate our model’s generalizability capabilities. I2P Schramowski et al. (2023b) consists of unsafe prompts only, which we combine with captions in the COCO 2017 validation set Lin et al. (2015) (assuming all captions are safe) to create the I2P++ dataset. We split I2P++ and UD into validation and test sets with a ratio of 30/70. The sizes of the source datasets are summarized in Table 1.
@@ 4.3 Retrieval Library Construction
To evaluate the model’s adaptability to external datasets, we constructed an external library using the I2P++ and UD datasets. We applied K-Means clustering to the safe and unsafe examples in these datasets, resulting in a total of 991 safe examples and 700 unsafe examples collected from the I2P++ and UD validation sets.
@@
To investigate the impact of library size on model performance, we created a series of smaller external libraries by downsampling the initial external library. Specifically, we generated three smaller libraries with sizes corresponding to 1/8, 1/4, and 1/2 of the original library size (Table 2). We achieved this downsampling by using K-Means to further cluster the full-size external library into smaller groups and selecting the centroid examples.
@@ 5 Experiments
We conducted a comprehensive experimental evaluation to assess the performance of our proposed model. To provide a thorough comparison, we selected two baseline models: WPIE (a 4-layer XLM-R) and LLAMA3 (Llama-3-8b), with the latter configured according to the settings outlined in Llama Guard Inan et al. (2023). Our experimental evaluation consisted of seven distinct components, which are detailed in the following sections.
The experimental setup is described in Section 5.1. We then present the results of our evaluation, which examined six key aspects of our model’s performance: (1) classification performance and robustness to adversarial attacks (Section 5.2); (2) adaptability to external data sources (Section 5.3); (3) ability to follow instructions (Section 5.4); (4) scalability of performance with respect to retrieval library size (Section 5.5); (5) impact of reference example numbers on performance (Section 5.6); and (6) effect of different embedding models on performance (Section 5.7).
@@ 5.1 Experimental Setup
During training, our models learned to assess the input text by leveraging retrieved reference examples. We employed a modified Chain-of-Thought (CoT) Wei et al. (2023) approach. CoT has been shown to improve the response quality of large language models. In contrast to the typical CoT setup, where answers are derived by the reasoning process, we opted to place the answer before the reasoning process to minimize latency. Specifically, we enforced the first token to be the answer, followed by a citation and a reasoning section (Figure 4). The citation indicates which reference examples were used to inform the assessment, while the reasoning section provides an explanation for the induced assessment. At evaluation time, we only output a single token and use the probability of the "unsafe" token as the unsafe probability.
@@ 5.2 Classification and Robustness
We conducted a comprehensive evaluation of Class-RAG, comparing its performance to two baseline models, WPIE and LLAMA3, on the CoPro in-distribution (ID) test set and out-of-distribution (OOD) test set. To assess the robustness of our model against adversarial attacks, we augmented the test sets with 8 common obfuscation techniques using the Augly library Papakipos and Bitton (2022). The results, presented in Table 3, demonstrate that Class-RAG outperforms both baseline models. Notably, both LLAMA3 and Class-RAG achieved an AUPRC score of 1 on the in-distribution and out-of-distribution test sets, indicating excellent classification performance. However, Class-RAG (DRAGON RoBERTa) exhibits superior robustness to LLAMA3 against adversarial attacks, highlighting its ability to maintain performance in the presence of obfuscated inputs.
@@ 5.3 Adaptability to External Data
To assess the impact of library size on performance, we also created downscaled versions of the external library, denoted as EX ($\frac{1}{8}$), EX ($\frac{1}{4}$), and EX ($\frac{1}{2}$), which were constructed by downsampling the full external library to $\frac{1}{8}$, $\frac{1}{4}$, and $\frac{1}{2}$ of its original size, respectively. Notably, this approach ensures that there is no label leakage.
Our results in Table 4 demonstrate that Class-RAG struggles to perform well on the I2P++ dataset when relying solely on the ID library, achieving an AUPRC score of only 0.229. However, by incorporating new reference examples from the full external library, we observe a substantial 245% improvement in AUPRC, reaching a score of 0.791. Furthermore, the model’s performance against adversarial attacks also improves significantly, with a relative increase of 188% from 0.235 to 0.677. We observe similar improvements on the UD dataset, where the AUPRC score increases from 0.917 to 0.985, and the performance against adversarial attacks improves from 0.914 to 0.976.
Table 4: AUPRC scores for Class-RAG on the I2P++ and UD external datasets, using various retrieval libraries. Higher AUPRC scores indicate better performance.
Obfuscations ID Lib ID +EX(1/8) Lib ID +EX(1/4) Lib ID +EX(1/2) Lib ID +EX Lib
@@
None 0.229 0.548 0.634 0.685 0.791
change_case 0.311 0.650 0.721 0.761 0.843
insert_punctuation_chars 0.183 0.240 0.254 0.273 0.318
insert_text 0.270 0.603 0.685 0.724 0.816
whitespace_chars 0.249 0.497 0.470 0.477 0.601
merge_words 0.261 0.599 0.689 0.723 0.815
replace_similar_chars 0.165 0.355 0.384 0.436 0.549
simulate_typos 0.211 0.527 0.621 0.630 0.742
split_words 0.234 0.495 0.525 0.486 0.613
AVERAGE 0.235 0.501 0.554 0.577 0.677
@@
None 0.917 0.966 0.973 0.977 0.985
change_case 0.937 0.978 0.983 0.985 0.991
insert_punctuation_chars 0.894 0.915 0.923 0.914 0.931
insert_text 0.924 0.970 0.976 0.981 0.988
whitespace_chars 0.925 0.965 0.960 0.959 0.971
merge_words 0.933 0.975 0.980 0.984 0.990
replace_similar_chars 0.864 0.927 0.933 0.942 0.953
simulate_typos 0.911 0.971 0.975 0.975 0.984
split_words 0.918 0.961 0.964 0.959 0.972
AVERAGE 0.914 0.959 0.963 0.964 0.974

### v2

# Class-RAG: Real-Time Content Moderation with Retrieval Augmented Generation

@@ Abstract
Robust content moderation classifiers are essential for the safety of Generative AI systems. In this task, differences between safe and unsafe inputs are often extremely subtle, making it difficult for classifiers (and indeed, even humans) to properly distinguish violating vs. benign samples without context or explanation. Scaling risk discovery and mitigation through continuous model fine-tuning is also slow, challenging and costly, preventing developers from being able to respond quickly and effectively to emergent harms. We propose a Classification approach employing Retrieval-Augmented Generation (Class-RAG). Class-RAG extends the capability of its base LLM through access to a retrieval library which can be dynamically updated to enable semantic hotfixing for immediate, flexible risk mitigation. Compared to model fine-tuning, Class-RAG demonstrates flexibility and transparency in decision-making, outperforms on classification and is more robust against adversarial attack, as evidenced by empirical studies. Our findings also suggest that Class-RAG performance scales with retrieval library size, indicating that increasing the library size is a viable and low-cost approach to improve content moderation.
@@ 1 Introduction
Recent advances in Generative AI technology have enabled new generations of product applications, such as text generation OpenAI (2023); Anthropic (2023); Dubey (2024), text-to-image generation Ramesh et al. (2021); Dai et al. (2023); Rombach et al. (2022), and text-to-video generation Meta (2024). Consequently, the pace of model development must be matched by the development of safety systems which are properly equipped to mitigate novel harms, ensuring the system’s overall integrity and preventing the use of Generative AI products from being exploited by bad actors to disseminate misinformation, glorify violence, and proliferate sexual content Foundation (2023).
To achieve this goal, traditional model fine-tuning approaches are often employed, with classifiers learning patterns from labeled content moderation text data leveraged as guardrails OpenAI (2023). However, there are many challenges associated with automating content moderation with fine-tuning. First, content moderation is a highly subjective task, meaning that inter-annotator agreement in labeled data is low, due to different interpretations of policy guidelines, especially on borderline cases Markov et al. (2023). Second, it is impossible to enforce a universal taxonomy of harm, not only due to the subjectivity of the task, but due to the impact of systems scaling to new locales, new audiences, and new use cases, with different guidelines and different gradients of harm defined on those guidelines Shen et al. (2024). Third, the fine-tuning development cycle, which encompasses data collection, annotation, and model experimentation, is not ideally suited to the content moderation domain, where mitigations must land as quickly as possible once vulnerabilities are established.
To address these challenges of subjectivity and inflexibility as a result of scale, we propose a Classification approach to content moderation which employs Retrieval-Augmented Generation (Class-RAG) to add context to elicit reasoning for content classification. While RAG Lewis et al. (2020) is often used for knowledge-intensive tasks where factual citation is key, we find that a RAG-based solution offers a distinct value proposition for the classification task of content moderation, not only due to its ability to enhance accuracy with few-shot learning, but because of its ability to make real-time knowledge updates, which is critical in our domain for speedy mitigations.
Our content moderation system consists of an embedding model, a retrieval library consisting of both negative and positive examples, a retrieval module, and a fine-tuned LLM classifier. When a user inputs a query, we retrieve the most similar negative and positive examples, and enrich the original input query to the classifier with the contextual information derived from similar retrieved queries.
@@
  • Real-time Mitigation: Class-RAG enables swift mitigation of generated content through its easily updated retrieval library, allowing changes to take effect within minutes to hours, contingent on retrieval library indexing speed. This approach significantly outpaces traditional model retraining, which typically requires several days to weeks.
  • Improved Classification Performance: Our experiments demonstrate that Class-RAG achieves superior classification performance compared to fine-tuning a lightweight 4-layer Transformer pre-trained on content moderation data and fine-tuning a general-purpose 8b parameter LLM.
  • Low-Cost Customization: By customizing the retrieval library, Class-RAG facilitates low-cost adaptation to diverse applications, allowing seamless policy updates without requiring model retraining. Maintaining multiple retrieval libraries is more cost-effective than building multiple models, reducing development, serving, and maintenance costs.
@@ 2 Related Work
Content moderation and Generative AI safety
Much work has been done in the last decade to mitigate the dissemination of undesired content in the wake of innovations in communication technologies. Machine learning approaches have been proposed to address sentiment classification Yu et al. (2017), harassment Yin et al. (2009), hate speech detection Gambäck and Sikdar (2017), abusive language Nobata et al. (2016), and toxicity Adams et al. (2017). General improvements in deep learning have also accelerated the field of content moderation. WPIE, or Whole Post Integrity Embeddings, built with BERT and XLM on top of advances in self-supervision, obtains a holistic understanding of a post through a pretrained universal representation of content Schroepfer (2019). Advances in Generative AI have also spurred the question of whether or not LLMs could potentially be used as content moderators Huang (2024). However, the capabilities of Generative AI introduce a proliferation of harm types beyond hate speech or toxicity detection whose mitigations and benchmarks engage further research and exploration. A comprehensive AI harm taxonomy encompasses such harm categories like academic dishonesty, unauthorized privacy violations, and non-consensual nudity Zeng et al. (2024). Studies establish the difficulty of moderating text-predictive models, finding that neural classifiers have stronger performance but occasionally unacceptable leakage (stronger precision) while extensive blocklists are more effective in harm mitigation but lead to unnecessary suppression (stronger recall) Vashishtha et al. (2023). OpenAI partially mitigates the weaknesses of neural classifiers by investing in data quality management and active learning Markov et al. (2023). Benchmarks establish baselines for the efficacy of existing classifiers and have provided valuable datasets to evaluate harmful categories like self-harm, illegal activity, sexual content, and graphic violence, such as UnsafeBench Qu et al. (2024), I2P  Schramowski et al. (2023a), and P4D Chin et al. (2024).
@@
Retrieval Augmented Generation (RAG) Lewis et al. (2020) improves the base capabilities of large pre-trained language models with a retrieval mechanism to explicit non-parametric memory, and has been demonstrated to mitigate problems with LLM outputs such as training cut-off, interpretability, and hallucination Zhao et al. (2024), showing particular success with knowledge-intensive tasks Gao et al. (2024). The flexibility of RAG-based approaches allows for applications that do not require additional in-domain finetuning. For example, RAFT improves the model’s ability to answer questions in open-book in-domain settings Zhang et al. (2024). The baseline capabilities of RAG can also be augmented by innovating on its components, such as retrieval, by improving the documents or embedding model. Employing LLM generations in conjunction with vanilla retrieval often results in better performance, potentially due to better utilization of the world knowledge stored in parameters Yu et al. (2023) or better identification of neighborhoods in the corpus embedding space Gao et al. (2022). DRAGON demonstrates that a fairly small BERT-based model can be trained for good performance on dense retrieval with an ensemble of data augmentation with diverse relevance labels Lin et al. (2023).
@@ 3 System Architecture
Figure 1: Architecture of Class-RAG. For comparison, Llama Guard is depicted without a retrieval model.
Class-RAG is a four-part system consisting of an embedding model, a retrieval library, a retrieval module, and a fine-tuned LLM classifier (Figure 1). When a user inputs a prompt, an embedding is computed on the prompt via the embedding model, which is compared against an index of embeddings for positive and negative prompts in the retrieval library. Using Faiss, a library for efficient similarity search Douze et al. (2024), $k$ nearest reference examples are retrieved against the embedding of the user input prompt, and the reference examples and input prompt are then sent to the fine-tuned LLM for classification.
@@ 3.1 Embedding Model
We leverage the DRAGON RoBERTa Lin et al. (2023) context encoder as our primary embedding model. DRAGON is a bi-encoder dense retrieval model utilizing a dual-encoder architecture to embed queries and documents into dense vector representations, facilitating efficient retrieval of relevant information. Ablations on embedding model are discussed in the Experiments section below.
@@ 3.2 Retrieval Library
Our retrieval library is comprised of two distinct sub-libraries: a safe library and an unsafe library. Each entry in the retrieval library is represented by four attributes: (1) prompt, (2) label, (3) embedding, and (4) explanation. The construction of the retrieval library is described in detail in the Data Preparation section.
@@ 3.2.1 Retrieval Module
Given the selected embedding, we leverage the Faiss library for similarity search Douze et al. (2024) to efficiently retrieve the two nearest safe and unsafe examples from the retrieval library, computing the L2 distance metric to establish the similarity between the input embedding and the embeddings stored in the retrieval library.
@@ 3.3 LLM Classifier
Inspired by Llama Guard Inan et al. (2023), the classifier is fine-tuned on top of the OSS Llama-3-8b checkpoint Dubey (2024). We leverage the CoPro dataset Liu et al. (2024) to train and evaluate our model.
@@ 4.1 Dataset Details
We leverage the CoPro dataset Liu et al. (2024) to train and evaluate our classifier. In addition to CoPro, we use the Unsafe Diffusion (UD) Qu et al. (2023) and I2P++ Liu et al. (2024) datasets to evaluate our model’s generalization capabilities. I2P Schramowski et al. (2023b) consists of unsafe prompts only, which we combine with captions in the COCO 2017 validation set Lin et al. (2015) (assuming all captions are safe) to create the I2P++ dataset. We split I2P++ and UD into validation and test sets with a ratio of 30/70. The sizes of the source datasets are summarized in Table 1.
@@ 4.3 Retrieval Library Construction
To assess the model’s adaptability to external datasets, we created an external library using the I2P++ and UD datasets. We applied K-Means clustering to the safe and unsafe examples in these datasets, with K set to 1000. After discarding clusters with fewer than 2 examples, our library consisted of 991 safe examples and 700 unsafe examples collected from the I2P++ and UD validation sets.
@@
To investigate the impact of library size on model performance, we generated a series of smaller external libraries by downsampling the original external library. Specifically, we created three smaller libraries, each containing 1/8, 1/4, and 1/2 of the external library’s examples (Table 2). To downsample, we reapplied K-Means clustering to the safe and unsafe examples in the full-size library, using a reduced number of clusters (K) proportional to the desired library size. For instance, for the EX(1/2) library, we set K to 500 (approximately half of 991) for safe examples and 350 (half of 700) for unsafe examples.
@@ 5 Experiments
We conducted a comprehensive experimental evaluation to assess the performance of our proposed model. To provide a thorough comparison, we selected two baseline models: WPIE (a 4-layer XLM-R) and LLAMA3 (Llama-3-8b), with the latter configured according to the settings outlined in Llama Guard Inan et al. (2023). Our experimental content consisted of seven distinct components, which are detailed in the following sections.
The experimental setup is described in Section 5.1. We then present the results of our evaluation, which examined six key aspects of our model’s performance: (1) classification performance and robustness to adversarial attacks (Section 5.2); (2) adaptability to external data sources (Section 5.3); (3) ability to follow instructions (Section 5.4); (4) impact of retrieval library size on performance (Section 5.5); (5) impact of reference example numbers on performance (Section 5.6); and (6) impact of embedding models on performance (Section 5.7).
@@ 5.1 Experimental Setup
During training, our models learned to assess the input text by leveraging retrieved reference examples. We employed a modified Chain-of-Thought (CoT) Wei et al. (2023) approach. CoT has been shown to improve the response quality of large language models. In contrast to the typical CoT setup, where answers are derived by the reasoning process, we opted to place the answer before the reasoning process to minimize inference latency. Specifically, we enforced the first token to be the answer, followed by a citation and a reasoning section (Figure 5). The citation indicates which reference examples were used to inform the assessment, while the reasoning section provides an explanation for the induced assessment. At inference time, we only output a single token and use the probability of the "unsafe" token as the unsafe probability.
@@ 5.2 Classification and Robustness
We conducted a comprehensive evaluation of Class-RAG, comparing its performance to two baseline models, WPIE and LLAMA3, on the CoPro test set. To assess the robustness of our model against adversarial attacks, we augmented the test sets with 8 common obfuscation techniques using the Augly library Papakipos and Bitton (2022). The results, presented in Table 3, demonstrate that Class-RAG outperforms both baseline models. Notably, both LLAMA3 and Class-RAG achieved an AUPRC score of 1 on the test set, indicating excellent classification performance. However, Class-RAG (DRAGON RoBERTa) exhibits superior robustness to LLAMA3 against adversarial attacks, highlighting its ability to maintain performance in the presence of obfuscated inputs.
@@ 5.3 Adaptability to External Data
As shown in Table 4, models trained on the CoPro dataset struggle to generalize to out-of-distribution external datasets, such as I2P++. In contrast, performance on the UD evaluation set is stronger, likely due to the similar distribution between UD and CoPro. Notably, Class-RAG’s performance on I2P++ is poor when relying solely on the in-distribution (ID) library, with an AUPRC score of only 0.229. However, incorporating new reference examples from the full external library leads to a substantial improvement in AUPRC, with a 245% increase to 0.791. This enhancement also translates to improved robustness against adversarial attacks, with a relative increase of 188% from 0.235 to 0.677. Similar improvements are observed on the UD dataset, where the AUPRC score rises from 0.917 to 0.985, and performance against adversarial attacks improves from 0.914 to 0.976.
Table 4: AUPRC scores on the I2P++ and UD external datasets. Higher AUPRC scores indicate better performance.
Obfuscations WPIE LLAMA3 Class-RAG (ID Lib) Class-RAG (ID+EX Lib)
@@
None 0.361 0.165 0.229 0.791
change_case 0.247 0.098 0.311 0.843
insert_punctuation_chars 0.171 0.114 0.183 0.318
insert_text 0.307 0.170 0.270 0.816
whitespace_chars 0.158 0.134 0.249 0.601
merge_words 0.289 0.166 0.261 0.815
replace_similar_chars 0.136 0.133 0.165 0.549
simulate_typos 0.180 0.145 0.211 0.742
split_words 0.142 0.140 0.234 0.613
AVERAGE 0.221 0.141 0.235 0.912
@@
None 0.949 0.867 0.917 0.985
change_case 0.917 0.671 0.937 0.991
insert_punctuation_chars 0.783 0.807 0.894 0.931
insert_text 0.938 0.844 0.924 0.988
whitespace_chars 0.860 0.792 0.925 0.971
merge_words 0.930 0.856 0.933 0.990
replace_similar_chars 0.817 0.750 0.864 0.953
simulate_typos 0.884 0.819 0.911 0.984
split_words 0.839 0.825 0.918 0.972
AVERAGE 0.880 0.803 0.914 0.974